OpenStudy (anonymous):

PLEASE HELP ME. Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. Provide the solution to the system. 5x – 3y = 1 7x – 4y = 2

OpenStudy (turingtest):

Let's number our equations: 5x – 3y = 1 (1) 7x – 4y = 2 (2) First the y seems easiest to eliminate; multiply (1) by 4 and (2) by -3 to get 20x – 12y = 4 (1) -21x + 12y = -6 (2) Now add (1) and (2) and the y's will cancel: -x=-2 --------> x=2 now that we know the value of x we can use either equation to solve for y, let's use (1) y=(1/3)(5x-1)=(5(2)-1)/3=3 so x=2 y=3 check your work by plugging these values into the original system and see if they yield true statements.

OpenStudy (anonymous):

thnks for your help...I have one question..you multiply equation (1) by 4 and equation (2) by -3..shouldnt it be -4 for (1) ??

OpenStudy (turingtest):

Not if you're going to ADD (1) to (2) to eliminate the y's. If you multiply (1) by -4 then you have to SUBTRACT it from (2) to eliminate the y's. I like to only add equations when possible, it helps me keep track of all the minus signs. When you subtract a whole equation it is like adding its negative, so when multiplying (1) by -4 and then subtracting (2) it seems to over-complicate the situation and it is more likely you will lose track of a minus sign somewhere. I looked for another question on your page but I just saw this one posted twice.

OpenStudy (anonymous):

thank you sooo much for your help